# On the Presence of Angular-Velocity Offsets in Disk Galaxies

### v1

@@ 5 Summary and Discussion
The generality of the property was analysed by comparing it to well-established models. Our analysis seem to support the presence of offsets in disk galaxies, as the property performs equally well (or better) than the standard models. Compared to a Burkert profile, it is preferred in 60% of the cases, while relative to NFW it is superior in 73% of the cases. The analysis was carried out by applying the SPARC sample, which consists of 175 rotation curves. As a potential next step, it would be beneficial to test the hypotheses on larger samples, like MaNGA (Wake et al., 2017).
A beneficial side-effect of the offsets is the introduction of a method that estimates galaxies’ total baryonic masses. By focusing on the peripheral regions, where a Keplerian behaviour is expected, we were able to design a velocity-model without assuming how the mass is distributed. The total baryonic mass is an explicit parameter of the new model while the parameters that define the distribution of the dark (or the baryonic) profiles are not necessary. Potentially, this method could be applied in the future to estimate the baryonic masses of galaxies. Future routes for investigation, in this context, may include the compatibility of this method with a baryonic Tully-Fisher relation (McGaugh et al., 2000), or the sensitivity of the masses to the degeneracy between the disk and the halo (Dutton et al., 2005).
@@
A possible mechanism of such a kind may be linked to a rotation of the reference frame (Gomel & Zimmerman, 2021). More accurately, a rotation of the local inertial frame where the model is applicable, relative to the frame where the observations are given. Such a rotation would induce an additional constant angular-velocity term in the data.
@@
The coupling between baryonic and dark matter in galaxies is a fundamental aspect of cosmic structure formation (Petit, Q. et al., 2023; Zhu et al., 2017). Observational studies, particularly those analysing galaxy rotation curves, have revealed that the distribution of baryonic matter can significantly influence the morphology of dark matter halos. For instance, high surface brightness galaxies exhibit more contracted and cuspy dark matter profiles, suggesting a strong coupling between baryonic distribution and halo shape (Li, Pengfei et al., 2022). Similarly, analyses employing Einasto halo fits have revealed correlations between the core properties of dark matter halos and the extent of the baryonic component (Ghari et al., 2019). The contribution of the current study is the introduction of an explicit relationship between the distributions.
The emergence of such an explicit relation may benefit from further exploration. One future possibility is to analyse the distribution by comparing it to standard dark halo profiles. Since each profile has its own distinctive characteristics, it may be illuminating to visualize the different profiles together. This may be particularly relevant from the perspective of the core-cusp problem (de Blok, 2010). Let us clarify that a deeper understanding of Eq. 11 is equivalent to understanding the offsets.
@@ Acknowledgements
This research was supported in part by the Israel Science Foundation (ISF), grant No. 1404/22.
@@
The rotation-curve data underlying this study are available in the DATA SECTION, at http://astroweb.cwru.edu/SPARC/. It includes extended 21-cm rotation curves and B-band photometry for the sample of galaxies analysed in this work. The complete set of figures supporting the findings of this study is openly available at https://doi.org/10.5281/zenodo.15182796. These include figures and tabulated data of the model’s fitting results.
Appendix A The Dark-Halo Density Profile
@@
Appendix B Angular-Velocity Curves
Fig. 8 presents a selection of angular-velocity curves, to provide the reader with a better sense of their behaviour, features, and relationship to baryonic components.

### v2

@@ 5 Summary and Discussion
The generality of the property was analysed by comparing it to well-established models. Our analysis seem to support the presence of offsets in disk galaxies, as the property performs equally well (or better) than the standard models. Compared to a Burkert profile, it is preferred in 60% of the cases, relative to NFW it is superior in 73% of the cases, and relative to MOND it is equally credible. The analysis was carried out by applying the SPARC sample, which consists of 175 rotation curves. As a potential next step, it would be beneficial to test the hypotheses on larger samples, like MaNGA (Wake et al., 2017).
In adopting the current property, a concern may emerge. A constant angular velocity term is equivalent to a circular velocity that increases linearly with radius. At a sufficiently large distance, the linear term may become dominant and lead to increasing rotation curves in the outer parts of galaxies. Yet, most RCs are observed to flatten out in these regions (Sofue and Rubin, 2001).
The resolution of this concern is fairly straightforward. A visual inspection of the full sample reveals that in most RCs the rate at which the baryonic term is declining is approximately the same as the rate at which the linear term rises over very long distances. This behaviour produces flat, slightly rising, or slightly falling curves at the outer parts of galaxies. In Appendix D we present this concern and its resolution in more detail, as well as a few examples.
A beneficial side-effect of the offsets is the introduction of a method that estimates galaxies’ total baryonic masses. By focusing on the peripheral regions, where the baryonic component exhibits a Keplerian behaviour, it was possible to design an analytical model for the rotational velocities. The proposed model for the velocities does not assume how the mass is distributed. The parameters that define the distribution of the dark (or the baryonic) profiles are not necessary. The only model-parameters are the total baryonic mass and the offset. This enables the calculation of baryonic masses while not being affected by baryonic modeling. Potentially, this method could be applied in the future to a large number of galaxies to estimate their baryonic masses.
Other routes for investigation, in this context, may include the compatibility of the masses with a baryonic Tully-Fisher relation (McGaugh et al., 2000). Since the outer flat parts of rotation curves can be well-fitted with the current model (see examples in Appendix D), it may be interesting to correlate the baryonic masses with their corresponding typical flat velocities, assuming the latter represent the Keplerian region. Another direction would be the sensitivity of the current method to the degeneracy between the disk and the halo (Dutton et al., 2005).
@@
A possible mechanism of such a kind may be linked to a rotation of the reference frame (Gomel and Zimmerman, 2021). More accurately, a rotation of the local inertial frame where the model is applicable, relative to the frame where the observations are given. Such a rotation would induce an additional constant angular-velocity term in the data.
@@
The coupling between baryonic and dark matter in galaxies is a fundamental aspect of cosmic structure formation (Petit, Q. et al., 2023; Zhu et al., 2017). Observational studies, particularly those analysing galaxy rotation curves, have revealed that the distribution of baryonic matter can significantly influence the morphology of dark matter halos. For instance, high surface brightness galaxies exhibit more contracted and cuspy dark matter profiles, suggesting a strong coupling between baryonic distribution and halo shape (Li, Pengfei et al., 2022). Similarly, analyses employing Einasto halo fits have revealed correlations between the core properties of dark matter halos and the extent of the baryonic component (Ghari et al., 2019). Since the current study introduces an explicit relationship between the two distributions, several observational studies in small-scale cosmology (e.g., Benisty et al. (2024)) may benefit from using it directly.
The emergence of an explicit relation may require further exploration. One future possibility is to analyse the relation by comparing it to standard dark halo profiles. Since each profile has its own distinctive characteristics, it may be illuminating to visualize the different profiles together as in Fig. 6. This may be particularly relevant from the perspective of the core-cusp problem (de Blok, 2010). Let us clarify that a deeper understanding of Eq. 11 is equivalent to understanding the offsets.
Another subject to consider is the connection between the current phenomenology and MOND. Although MOND is based on theoretical assumptions, its application to RCs may be considered semi-empirical. This is due to the flexibility afforded by the selection of the unknown function $\mu(a/a_{0})$. In that sense, both MOND and the current model are semi-empirical prescriptions for rotation-curve fitting based on the corresponding baryonic component. In fact, both methods are highly sensitive to baryonic modeling. The current model uses a constant offset, which preserves any features in the baryonic term, while MOND uses only the baryonic term, together with the chosen $\mu(a/a_{0})$ (Milgrom, 1983).
There are also notable differences between the two models. First, for a given MOND function $\mu$ and a global acceleration value $a_{0}$, MOND requires only one free parameter per galaxy, that is, $\rm M/L$. The current model requires both $\rm M/L$ and an offset for each galaxy (see Appendix A). The second difference is more basic. Because MOND is a theory of gravity, acceleration is the most crucial quantity. The present phenomenology, however, suggests that angular velocities are the important metric to investigate.
@@ Acknowledgements
The authors wish to thank David Benisty for his helpful suggestions and the anonymous referee for his thoughtful questions and valuable feedback. This research was supported in part by the Israel Science Foundation (ISF), grant No. 1404/22.
@@
The rotation-curve data underlying this study are available in the DATA SECTION, at http://astroweb.cwru.edu/SPARC/. It includes extended 21-cm rotation curves and B-band photometry for the sample of galaxies analysed in this work. The complete set of figures supporting the findings of this study is openly available at https://doi.org/10.5281/zenodo.17392323. These include figures and tabulated data of the model’s fitting results.
Appendix A A comparison with MOND
One of the main goals of the present study is to establish the universality of constant offsets. In Section 2 we compare the current phenomenological model (i.e., of constant offsets) with existing models. A comparison with MOND, one of the earliest and most established models, may require several implementations. Therefore, the current appendix describes several different MOND versions and their corresponding results.
First, we follow Wang and Chen (2021) and choose a simple interpolation function of the form:
$\rm\mu(x)=\frac{x}{1+x},$ (12)
where $\mu$ is the MOND interpolation function, and $\rm x=a/a_{0}$, is the ratio between actual accelerations and the critical acceleration. For the critical acceleration we adopt a value of $\rm a_{0}=1.2\cdot 10^{-10}[m/s^{2}]$ (Begeman et al., 1991; Li, Pengfei et al., 2018).
To calculate MOND’s rotational velocities (i.e., to produce a MOND model), one needs a conversion formula translating from Newtonian velocities to MOND. The current formula can be directly derived from Eq. 12 and is explicitly given in Wang and Chen (2021).
Next, with a MOND model in hand, we apply an MCMC simulation to fit the model to the data. We use a single parameter (the stellar disk $\rm M/L$) with a positive prior ($\rm M/L>0$) to find the best-fit value for each galaxy and its corresponding $\chi^{2}$. The rest of the MCMC details are the same as in Section 2.
The MOND MCMC results are then used for a statistical comparison. We follow the scheme described in Section 2 to calculate the $\rm\Delta BICs$ between MOND and the current model. The results are shown in the left panel of Fig. 7. The $\rm\Delta BIC$ test already accounts for the fact that MOND requires only one free parameter.
According to the test, the present phenomenological model is preferred over MOND. The median value is 8.31 $\pm$ 1.74 with $65\%$ of the galaxies produce $\rm\Delta BIC>2$. Flawed results were excluded from the test (43 galaxies where MOND fails, 6 where the phenomenological model fails, and 30 where both models fail).
The next step is applying a different interpolation function. We rely on Li, Pengfei et al. (2018) and apply the following formula to convert the accelerations:
$\rm a=\frac{a_{bar}}{1-e^{-\sqrt{\frac{a_{bar}}{a_{0}}}}},$ (13)
where $\rm a$ is the MOND acceleration and $\rm a_{bar}$ is the baryonic Newtonian acceleration. The conversion in Eq. 13 is based on the radial-acceleration relation (RAR).
We then apply precisely the same scheme (MCMC, statistical analysis) to compare MOND and the current model. The results are shown in the central panel of Fig. 7. The histogram is almost unaffected by the two different MOND models. In fact, the vast majority of MOND fits are almost indistinguishable when visually inspected. In this case, the present phenomenological model is preferred over MOND, with a median value of 8.26 $\pm$ 1.73 and with $65\%$ of the galaxies producing $\rm\Delta BIC>2$. The excluded galaxies are similar to the previous test.
The last step is to treat the distance and inclination as free parameters. Following Li, Pengfei et al. (2018) we allow the distance and inclination to be adjusted while priors are set according to their known uncertainties. We also use a prior for the $\rm M/L$ parameter, in accordance with stellar population synthesis models and precisely as in Li, Pengfei et al. (2018). The extra flexibility considerably improves MOND results.
To enable a proper comparison, it is necessary to apply these adjustments to the phenomenological model as well. The details of the later analysis (that is, of the adjusted phenomenological model) are presented in Appendix B. The results of the comparison are shown in the right panel of Fig. 7.
According to the last test, the present phenomenological model and MOND are equally credible. The median value is -0.19 $\pm$ 0.68, with $36\%$ of the galaxies produce $\rm\Delta BIC>2$ and $39\%$ produce $\rm\Delta BIC<-2$. Throughout the main text, this final test is presented. Flawed results were excluded from the test (22 galaxies where MOND fails, 15 where the phenomenological model fails, and 31 where both models fail).
Figure 7: Comparing the phenomenological model with MOND. Left panel: A histogram of $\rm\Delta BIC$ values, where the comparison is relative to a simple MOND (see the main text). The median value of the histogram is 8.31 $\pm$ 1.74. Central panel: A comparison relative to a RAR-MOND profile. The median value is 8.26 $\pm$ 1.73. Right panel: A comparison relative to RAR-MOND with adjustable distance and inclination. The median value is -0.19 $\pm$ 0.68.
Appendix B An additional Fitting Method
The current appendix offers an additional method to fit the phenomenological model to the data. Following previous studies, and especially Li, Pengfei et al. (2018), the distance and inclination of a galaxy may be treated as free parameters, with priors assigned according to their uncertainties. This affords some extra flexibility in the fitting process and reflects an inherent uncertainty in the modeling.
In our case, it would be worthwhile to investigate whether such flexibility could produce lower $\rm M/L$ values. As discussed in Section 2, the current $\rm M/L$ values, obtained from the baseline fitting, are somewhat higher than expected by stellar population synthesis models (Bell and de Jong, 2001; Meidt et al., 2012). Another motivation to implement the current method is to enable the third statistical test in Appendix A.
In order to implement the outlined method we follow Li, Pengfei et al. (2018) and adjust the baryonic component through:
$\rm v_{bar}^{{}^{\prime}}=\sqrt{\frac{D}{D_{0}}}\ {v_{bar}},$ (14)
where $\rm D$ is the adjustable distance and $\rm D_{0}$ is the default value of the distance as given in SPARC data tables. The added parameter $\rm D$ is not entirely flexible. It is constrained by a Gaussian prior with its standard deviation being equal to the specified distance uncertainty.
A second additional parameter is disk inclination (Li, Pengfei et al., 2018). When the disk inclination is adjusted, the observed velocities are transformed by:
$\rm v_{obs}^{{}^{\prime}}=\frac{sin(i_{0})}{sin(i)}\ {v_{obs}},$ (15)
where $\rm i$ is the adjustable inclination and $\rm i_{0}$ is the default value of the inclination as given in SPARC data tables. The inclination is also constrained by a Gaussian prior with its standard deviation equal to the specified inclination uncertainty.
For the $\rm M/L$ parameter we use a log-normal prior with a mean value of $\rm 0.5\ [M_{\odot}/L_{\odot}]$ and a standard deviation of $\rm 0.15$ dex, the latter being a slightly weaker than the commonly used $0.1$-dex prior (Meidt et al., 2012; Li et al., 2020). We then run the MCMC simulation (as specified in Section 2) to obtain the best-fit parameters and their corresponding $\chi^{2}$ values.
The fitting results are quite informative. The number of unsuccessful fittings (with p-value < 0.001) increased from 36 to 46. It reflects galaxies where the $\chi^{2}$ is not significantly improved by the extra flexibility, but the higher number of parameters reduces their p-value. A reduction in p-value may also occur in galaxies with excellent fittings, where extra flexibility is not required. On the other hand, a few fittings (e.g., UGC-00731) have been considerably improved, with inclination being the dominant parameter.
Fig. 8 draws an updated histogram of $\rm M/L$ values. The median value of log($\rm M/L$) is -0.17 with a standard error of 0.05. This is substantially lower than the values in Fig. 3, but still a bit higher than expectations (Meidt et al., 2012). Whether this reflects a genuine property of the $\rm M/L$ values or a weakness in the phenomenological model deserves further exploration.
A final notable point deals with inclination values. Several galaxies exhibit phenomenological-model inclinations that strongly disagree with those obtained from a similar MOND analysis (i.e. those in Li, Pengfei et al. (2018)). As an example, UGC-08837 is well fitted by both models, with the phenomenological model producing $\rm i/i_{0}=1$ and MOND predicting $\rm i/i_{0}=0.73$. Future examination of the observed inclinations may provide insight into this conflict.
Figure 8: Histogram of updated $\rm M/L$ values. The median value of log($\rm M/L$) is $\rm-0.17\pm 0.05$.
Appendix C The Dark-Halo Density Profile
@@
Appendix D Rotation Curves behaviour at large radii
Figure 9: Three examples of galactic rotation curves. Top panel: NGC-0801, central panel: Milky-Way, and bottom panel: NGC-0100.
The current study proposes that the difference between an observed AVC and its corresponding baryonic prediction is constant. When adopting this property, a concerning behaviour may arise. In the outer parts of a galaxy, the constant angular velocity term may become dominant, leading to a circular velocity that increases linearly with radius. However, most RCs tend to flatten out at large radii (Sofue and Rubin, 2001).
To address this apparent difficulty, it would be beneficial to examine a few examples. Fig. 9 shows three rotation curves, with their data extending to more than four effective radii. The top panel presents a very extended RC (NGC-0801), which demonstrates the model’s capability to support long flat behaviour. The best-fit parameters are: $\rm M/L=0.63\pm 0.02$ $\rm M_{\odot}/L_{\odot}$], $\omega_{0}=0.41\pm 0.02$ $\rm[10^{-16}\,rad/s]$. The central panel presents the declining RC of the Milky-Way, which demonstrates the model’s ability to support galaxies with a dominant baryonic contribution. The data were obtained from a recent study (Jiao et al., 2023). In this example, an analytical model with a Freeman disk (Freeman, 1970) and a Hernquist bulge (Hernquist, 1990) was employed. The best-fit parameters are: $\rm M_{d}=12.2\pm 1.2$ $\rm 10^{10}\,M_{\odot}]$, $\rm R_{d}=5.1\pm 0.3$ $\rm[Kpc]$, $\rm M_{b}=1.5\pm 0.3$ $\rm 10^{10}\,M_{\odot}]$, $\rm R_{b}=0.30\pm 0.05$ $\rm[Kpc]$ and $\omega_{0}=0.31\pm 0.12$ $\rm[10^{-16}\,rad/s]$. Priors for the bulge and disk scale lengths were set according to known values (Jiao et al., 2023). The bottom panel displays a moderately rising RC (NGC-0100), illustrating the model’s ability to support a monotonically increasing behaviour. The best-fit parameters are: $\rm M/L=0.75\pm 0.07$ $\rm M_{\odot}/L_{\odot}$], $\omega_{0}=1.80\pm 0.13$ $\rm[10^{-16}\,rad/s]$.
In the first two cases, the rate at which the baryonic term declines is approximately the same as the rate at which the linear term increases over long distances. A visual examination of the entire sample reveals that this kind of behaviour is relatively general. This means that the linear term does not necessarily dictate a linearly rising pattern in the outer parts of an RC. Generally speaking, the model can produce flat, slightly rising, or slightly falling curves in the outer regions of RCs, and the linear term would typically dominate only at very large distances, usually between 2-8 $R_{\rm max}$.
In addition, we remind the data-driven nature of the current research. The model introduced in Eq. 1 is motivated by its compatibility with rotation curve data. It does not provide a complete theoretical framework that predicts the velocities in any particular situation, e.g. as in Mistele et al. (2024). Future studies may reveal the extent to which this model should be considered valid.
Appendix E Angular-Velocity Curves
Fig. 10 presents a selection of angular-velocity curves. These plots are not fits, but rather a visual representations of the data together with the corresponding typical baryonic components. The $\rm M/Ls$ of the baryonic components were taken from Li et al. (2020). Plots of this type provided the initial motivation to investigating the universality of constant offsets.